OpenStudy (anonymous):

PLEASE HELP I am stuck- I have been waiting for help for a half an hour Use part I of the Fundamental Theorem of Calculus to find the derivative of the integral from 4x to 5x of (u+5/u-1)du I will also show this as an equation below

OpenStudy (anonymous):

\[g(x)=\int\limits_{4x}^{5x}\frac{ u+5 }{ u-1 }du\]

OpenStudy (anonymous):

I know that this can be separated into two different integrals \[\int\limits_{0}^{5x}\frac{ u+5 }{ ?u-1}du +\int\limits_{4x}^{0}\frac{ u+5 }{ ?u-1}du\]

OpenStudy (slaaibak):

Substituting u - 1 = k dk = du u + 5 = k + 6 Upper bound: 5x - 1 Lower bound: 4x -1 Can you do it from here?

OpenStudy (anonymous):

can you step me through it? I want to make sure I don't mess it up :)

OpenStudy (slaaibak):

Yeah sure. \[\int\limits_{4x -1 }^{5x -1} {k + 6 \over k} dk = \int\limits_{4x-1}^{5x-1} {1 + {6 \over k}}dk\]

OpenStudy (slaaibak):

so the anti derivative of that is k + 6ln|k| Now just evaluate at the points: 5x - 1 + 6ln|5x-1| - (4x - 1 + 6ln|4x-1|) = x + 6ln|5x-1| - 6ln|4x-1| \[x + 6\ln({5x-1 \over 4x-1})\]

OpenStudy (anonymous):

for some reason my online homework isn't recognizing that as the correct answer.

OpenStudy (slaaibak):

That's odd, I'm pretty sure that's correct tho.

OpenStudy (slaaibak):

Try it with the | | instead of the brackets

OpenStudy (anonymous):

I think it wants it not with the anti-derivative the first fundamental theorem I think doesn't take the anti-derivate the second theorem does.

OpenStudy (slaaibak):

Ohh, mybad. I see. So what you wrote at the start is correct. Just change it to: \[\int\limits_{0}^{5x} {u + 5 \over u - 1} du - \int\limits_{0}^{4x}{u + 5 \over u-1} du\]

OpenStudy (slaaibak):

Now I'll think of how to apply the FTC pt1. Did this a long while ago, so gonna take some time haha

OpenStudy (anonymous):

Thank you I appreciate it!

OpenStudy (slaaibak):

So according to the FTC1, you must find the F(x) function. Where F(x) is the anti derivative of f(x) Rewrite f(u) as \[{u + 5 \over u - 1 } = {u - 1 + 6 \over u - 1} = 1 + {6 \over u - 1}\] So f(u) = 1 + 6/(u-1) and F(u) = u + 6ln(u-1) Now F(5x) = 5x + 6ln(5x-1) and F(4x) = 4x + 6ln(4x-1) so F(5x) - F(4x) = x + 6ln(5x-1) - 6ln(4x-1) which is the same as the previous method.